Shift report 20010909 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Sep 09 2001 - 07:52:20 EDT

  • Next message: Apache: "Shift report 20010909 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Sep  9 07:52:20 2001, Shift: 00:00-08:00
    Supervisor: sjs, with: none
    Log book#8 pages 97 - 98
    
    Shift Summary:
    ------------------------------------------------------------
    1.) About one hour (00:15-1:20) of useful beam at 
    beginning of shift,then store lost with "an Aux Power 
    Supply needs to be replaced...".   At 05:10 beam restored 
    for about 5 minutes, when a "Pass dropped" occurred.  
    At end of shift "Pass personnel are in...no estimate 
    for beam"
    2.) Used limited beam to adjust vertex position (done
    during run 4645).  MCR made four 0.1mm shifts in
    "+direction blue horizontal plane".  Starting with
    <BBz>=22.7 cm, the <BBz> after the sequential shifts
    were 17.3 cm, 9.2 cm, 5.8 cm, -2.0 cm, respectively.
    The adjustment went smoothly and only took about 20
    minutes.  
    
    Started run 4646 with relocated intersection, but beam
    lost after about 30 minutes of running.  No additional
    useful running during shift.
    3.) The pedestal for tile2_6 seems to be drifting
    substantially, even during the short pedestal calibration
    run.  This was also seen last night and suggests a 
    problem with this detector.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Sep 09 2001 - 07:52:55 EDT